# Physics-informed machine learning for combustion: A review

### v1

@@ 8.2 Discussions on implementation choices
Forward vs inverse problems. As shown in Tables 4 and 5, PIML is currently used more for inverse problems than for forward problems, and this is not limited to the field of combustion. There are many reasons for this: (1) Inverse problems are usually simpler than forward problems when both types of problems can be studied because of the involvement of data. When solving forward problems fails, inverse problems are the only option. (2) Inverse problems are more common than forward problems in practice. Forward problems have strict requirements and need complete equations and IC/BCs, while inverse problems have more relaxed requirements. Therefore, it is not that researchers do not study forward problems, but that the specific problems themselves are inverse problems. (3) Inverse problems have richer application significance than forward problems in terms of field reconstruction, parameter inference, physical modeling, and kinetics discovery. Forward problems often only involve simulation, which can also be achieved by traditional numerical methods. (4) Inverse problems are better suited to leveraging the advantages of PIML than forward problems. It is well known that the overall performance of PIML in forward problems is far inferior to that of traditional numerical methods 434. However, one advantage of PIML is that it can easily couple data and incomplete equations to solve inverse problems, which is not straightforward for traditional numerical methods. In fact, the data-driven characteristic is an inherent advantage of ML, which can be leveraged in PIML. Nevertheless, the forward problems still have two important aspects: (1) In terms of academic research, the forward problem is the basis of the inverse problem. Research on the forward problems helps deepen our understanding of PIML, which in turn helps us understand the inverse problems. Furthermore, many proposed methods on the forward problems in terms of representation models, optimization processes, and PDE systems can be seamlessly transferred to the inverse problems. (2) In terms of practical applications, although PIML is inferior to traditional solvers in solving single PDE, it has demonstrated advantages in parametric/surrogate solving. Examples can be found in Section 2.3.4, Section 2.4, and Reference 301. Compared with traditional solvers, it is more efficient, provides a continuous representation in the parameter space, and can easily perform parameter sensitivity analysis. Note that parametric/surrogate modeling can also be implemented in the way of inverse problems, where the data can be utilized. To sum up, both types of problems are important. For researchers and engineers, the choice between forward problems and inverse problems should depend on the specific circumstances and objectives.
@@
Continuous vs discrete PINNs. As defined in Section 2.3.1, continuous PINNs follow the coordinate-to-variable mapping way, i.e., $\widehat{\mathbf{u}}=\mathrm{NN}(\mathbf{z})$, while discrete PINNs follow the field-to-field mapping way, i.e., $\widehat{\mathbf{U}}_{\mathrm{out}}=\mathrm{NN}(\mathbf{U}_{\mathrm{in}})$. Both have their own advantages and disadvantages 136. Continuous PINNs provide a continuous representation of the solution, making them mesh-independent and allowing for easy query of values at any points. However, this greatly increases the difficulty of learning, as it requires learning the entire solution in the spatiotemporal domain. In contrast, discrete PINNs only need to learn the field-to-field mapping. The AD used in continuous PINNs eliminates the error in differentiation, but also increases the computational cost. Additionally, continuity makes it difficult for continuous PINNs to learn large gradients and discontinuous features, which are easier for discrete PINNs. Moreover, discrete PINNs are easier to implement hard constraints on PDEs, which will be elaborated in the part of "soft vs hard constraints" in this section. The choice between continuous PINNs and discrete PINNs should depend on the specific task type. For example, continuous PINNs are more suitable for learning continuous spatiotemporal solutions, reconstructing non-grid-based distributed data, and performing zero-error differentiation. Discrete PINNs are more suitable for tasks such as surrogate modeling of chemical kinetics (Section 4.2), super-resolution (Section 5.5), and AR simulations 116.
@@
Complete/detailed vs reduced/simplified physics. As mentioned in Sections 8.1.1 and 8.1.2, not all studies on PIML for combustion adopt complete/detailed physical models, such as detailed reaction mechanisms, complete PDEs, non-averaged NS equations, and closed turbulence terms. The choice between complete/detailed or reduced/simplified physical models should depend on the specific circumstances. For forward problems, the physical model can be simplified but must be complete, otherwise the problem will be ill-posed. For inverse problems with data, the choice is more flexible, and either a complete/detailed or a reduced/simplified physical model can be used. In fact, sometimes using only a few physical constraints can greatly enhance the overall performance. For example, some studies in Sections 4.2 and 4.3.1 show that using only species and element conservations can achieve a significant improvement. When there are discrepancies between the physics, model, and data, a reduced/simplified physical model can be used. See Section 8.1.2 for details. When both complete models and data are available, comparative studies can be conducted to identify which models have minimal impact on the results and can be omitted, thereby reducing computational costs, as demonstrated in Reference 307.
@@
Soft vs hard constraints. Classical PIML integrates physical laws through loss functions, which is a soft constraint, so there is still a risk of violating the laws. This is why hard constraints are necessary. As mentioned in the Section 2.3.4, there are many ways to implement hard constraints for IC/BCs, PDEs, and other physical laws. For example, IC/BCs for continuous PINNs can be enforced through input and output transformations, while for discrete PINNs, they can be enforced through padding operations. Discrete PINNs can also easily implement hard constraints on PDEs using convolution kernels or traditional numerical methods/solvers. However, it is difficult to implement hard constraints on PDEs for continuous PINNs. Even for the hard constraint projection (HCP) 111, it is still necessary to minimize the PDE losses after projection. This is one of the advantages of discrete PINNs over continuous PINNs. For physical laws in the form of algebraic equations, i.e., most of the equations in Section 3, hard constraints can be implemented simply by directly specifying them, for both continuous and discrete PINNs. For example, the law of mass action and the Arrhenius law are hard constraints in all the literature that considers them introduced in this paper, because they are explicitly embedded in the intermediate steps of the method rather than constrained by the loss. Hard constraints have both advantages and disadvantages. Imposing hard constraints on certain physical laws will alleviate the burden of loss optimization. For example, enforcing species conservation can reduce one equation of $Y$ that need to be solved. However, hard constraints can sometimes increase the difficulty of optimization, especially for IC/BCs, as they often involve hyperparameters and "hyperfunctions". In some cases, such as PDEs, complexity is also a disadvantage of hard constraints, which are less straightforward than soft constraints. The choice between soft constraints and hard constraints depends on many factors. By default, hard constraints should be considered first. Soft constraints can be used in the following situations: (1) when the constraint requirements are not strict; (2) when hard constraints are difficult to implement, such as for PDEs in continuous PINNs; (3) when hard constraints are extremely cumbersome to implement; (4) when hard constraints impair or greatly slow down the optimization and convergence, which requires a posteriori testing.
@@ 8.3 Opportunities for combustion and AI
  • Academic research. First, there are still many limitations that need to be further solved, as mentioned in Section 8.1, including large-domain long-time problems, stiffness caused by detailed reaction mechanisms, model-data-physics discrepancies, and reproducibility issues. Second, there are still many combustion scenarios that can be studied using PIML, such as multiphase combustion, pollutant control, combustion diagnostics, microgravity combustion, explosions, and fires. Finally, there are still many tasks that can be achieved using PIML, such as improving combustion models, gaining insights from experimental data, and developing hybrid solvers.
@@
  • Integration of advanced AI methods. As introduced in this paper, various AI methods haven been used in PIML for combustion, including the basic MLP, KAN, NO, NODE, GAN, AE, and LSTM. However, there are still many advanced AI methods that have potential for application, such as DM, self-supervised learning, reinforcement learning, transfer learning, foundation models, and PDE discovery. Among these, DM has already been applied to fluid mechanics 166, 437, 438. Foundation models are a hot topic due to their notable applications in the NLP field, including LLMs such as GPT 24, Llama 439, and DeepSeek 440. In the SciML field, foundation models are a current key direction 86, 441. Some foundation models for PDEs have been developed, including the PDEformer 442, 443, UPS 444, Poseidon 445, PROSE-PDE 446, and PROSE-FD for fluid PDEs 447. Therefore, establishing foundation models for combustion is a feasible and promising direction, which also needs to be informed by combustion physics.
@@
  • Combustion-inspired AI. Physics-inspired AI paradigms are an important topic, in which AI methods are often proposed or improved by drawing inspiration from physical laws, phenomena, or principles. The inspiration can be achieved in many ways, including embedding physical laws, mimicking physical processes, and borrowing physical principles. Therefore, physics-inspired AI includes physics-informed AI, or it can be considered as physics-informed AI in a broad sense. Jiao et al. 448 have reviewed the AI methods that are inspired by classical mechanics, electromagnetism, statistical physics, and quantum mechanics. Additionally, two currently popular generative models are also inspired by physics: DM 160, 161 is inspired by nonequilibrium thermodynamics, while the continuity equation in the flow matching 449 originates from fluid mechanics. Therefore, it is natural to ask: can we develop combustion-inspired AI? In fact, CRNN can be regarded as one such effort 75, with its network architecture designed based on the law of mass action and Arrhenius law. However, given the abundance of combustion laws, phenomena, and principles, there are still many possibilities to explore, which may give rise to some revolutionary AI methods.
@@ 9 Conclusions
This paper provides a comprehensive review on PIML for combustion, systematically summarizing the its fundamental principles, major studies, key advances, and available resources, while also providing constructive discussions and recommendations for researchers and engineers. The main conclusions are as follows: (1) PIML has been widely and successfully applied to combustion chemical kinetics, combustion reacting flows, and other combustion problems, addressing both forward and inverse problems. These applications demonstrate the growing attention, proven effectiveness, and significant potential of PIML in combustion research. (2) Numerous approaches exist for integrating combustion laws into ML. A primary focus of this review is to explain how such laws can be integrated into ML frameworks–either through soft or hard constraints, via loss functions or representation models, within coordinate-to-variable or field-to-field paradigms, and using either complete/detailed or reduced/simplified physics–highlighting the flexibility of PIML. (3) By incorporating combustion knowledge, the physical consistency and accuracy of ML models are always enhanced compared to purely data-driven ML. (4) PIML serves as a unified bridge between physics, models, and data, enabling both model-to-data simulation/reconstruction tasks and data-to-model modeling tasks under the guidance of physics This integration produces three key outcomes that represent or contain physics: enhanced data, improved physical models, and more reliable ML models. When discrepancies arise among physics, physical models, and data, PIML can address them through a range of strategies. (5) Challenges remain in terms of the inherent limitations of PIML, physics-model-data discrepancies, and reproducibility. Furthermore, many potential applications and research topics remain unexplored. Both of these points indicate promising directions for the future. In conclusion, PIML for combustion is a rapidly evolving field that calls for increased collaboration between combustion and AI researchers to address significant problems, develop improved strategies, and explore emerging opportunities.

### v2

@@ 8.2 Discussions on implementation choices
Forward vs inverse problems. As shown in Tables 4 and 5, PIML is currently used more for inverse problems than for forward problems, and this is not limited to the field of combustion. There are many reasons for this: (1) Inverse problems are usually simpler than forward problems when both types of problems can be studied because of the involvement of data. When solving forward problems fails, inverse problems are the only option. (2) Inverse problems are more common than forward problems in practice. Forward problems have strict requirements and need complete equations and IC/BCs, while inverse problems have more relaxed requirements. Therefore, it is not that researchers do not study forward problems, but that the specific problems themselves are inverse problems. (3) Inverse problems have richer application significance than forward problems in terms of field reconstruction, parameter inference, physical modeling, and kinetics discovery. Forward problems often only involve simulation, which can also be achieved by traditional numerical methods. (4) Inverse problems are better suited to leveraging the advantages of PIML than forward problems. It is well known that the overall performance of PIML in forward problems is far inferior to that of traditional numerical methods 441. However, one advantage of PIML is that it can easily couple data and incomplete equations to solve inverse problems, which is not straightforward for traditional numerical methods. In fact, the data-driven characteristic is an inherent advantage of ML, which can be leveraged in PIML. Nevertheless, the forward problems still have two important aspects: (1) In terms of academic research, the forward problem is the basis of the inverse problem. Research on the forward problems helps deepen our understanding of PIML, which in turn helps us understand the inverse problems. Furthermore, many proposed methods on the forward problems in terms of representation models, optimization processes, and PDE systems can be seamlessly transferred to the inverse problems. (2) In terms of practical applications, although PIML is inferior to traditional solvers in solving single PDE, it has demonstrated advantages in parametric/surrogate solving. Examples can be found in Section 2.3.4, Section 2.4, and Reference 307. Compared with traditional solvers, it is more efficient, provides a continuous representation in the parameter space, and can easily perform parameter sensitivity analysis. Note that parametric/surrogate modeling can also be implemented in the way of inverse problems, where the data can be utilized. To sum up, both types of problems are important. For researchers and engineers, the choice between forward problems and inverse problems should depend on the specific circumstances and objectives.
@@
Continuous vs discrete PINNs. As defined in Section 2.3.1, continuous PINNs follow the coordinate-to-variable mapping way, i.e., $\widehat{\mathbf{u}}=\mathrm{NN}(\mathbf{z})$, while discrete PINNs follow the field-to-field mapping way, i.e., $\widehat{\mathbf{U}}_{\mathrm{out}}=\mathrm{NN}(\mathbf{U}_{\mathrm{in}})$. Both have their own advantages and disadvantages 139. Continuous PINNs provide a continuous representation of the solution, making them mesh-independent and allowing for easy query of values at any points. However, this greatly increases the difficulty of learning, as it requires learning the entire solution in the spatiotemporal domain. In contrast, discrete PINNs only need to learn the field-to-field mapping. The AD used in continuous PINNs eliminates the error in differentiation, but also increases the computational cost. Additionally, continuity makes it difficult for continuous PINNs to learn large gradients and discontinuous features, which are easier for discrete PINNs. Moreover, discrete PINNs are easier to implement hard constraints on PDEs, which will be elaborated in the part of "soft vs hard constraints" in this section. The choice between continuous PINNs and discrete PINNs should depend on the specific task type. For example, continuous PINNs are more suitable for learning continuous spatiotemporal solutions, reconstructing non-grid-based distributed data, and performing zero-error differentiation. Discrete PINNs are more suitable for tasks such as surrogate modeling of chemical kinetics (Section 4.2), super-resolution (Section 5.5), and AR simulations 119.
@@
Complete/detailed vs reduced/simplified physics. As mentioned in Sections 8.1.1 and 8.1.2, not all studies on PIML for combustion adopt complete/detailed physical models, such as detailed reaction mechanisms, complete PDEs, non-averaged NS equations, and closed turbulence terms. The choice between complete/detailed or reduced/simplified physical models should depend on the specific circumstances. For forward problems, the physical model can be simplified but must be complete, otherwise the problem will be ill-posed. For inverse problems with data, the choice is more flexible, and either a complete/detailed or a reduced/simplified physical model can be used. In fact, sometimes using only a few physical constraints can greatly enhance the overall performance. For example, some studies in Sections 4.2 and 4.3.1 show that using only species and element conservations can achieve a significant improvement. When there are discrepancies between the physics, model, and data, a reduced/simplified physical model can be used. See Section 8.1.2 for details. When both complete models and data are available, comparative studies can be conducted to identify which models have minimal impact on the results and can be omitted, thereby reducing computational costs, as demonstrated in Reference 313.
@@
Soft vs hard constraints. Classical PIML integrates physical laws through loss functions, which is a soft constraint, so there is still a risk of violating the laws. This is why hard constraints are necessary. As mentioned in the Section 2.3.4, there are many ways to implement hard constraints for IC/BCs, PDEs, and other physical laws. For example, IC/BCs for continuous PINNs can be enforced through input and output transformations, while for discrete PINNs, they can be enforced through padding operations. Discrete PINNs can also easily implement hard constraints on PDEs using convolution kernels or traditional numerical methods/solvers. However, it is difficult to implement hard constraints on PDEs for continuous PINNs. Even for the hard constraint projection (HCP) 114, it is still necessary to minimize the PDE losses after projection. This is one of the advantages of discrete PINNs over continuous PINNs. For physical laws in the form of algebraic equations, i.e., most of the equations in Section 3, hard constraints can be implemented simply by directly specifying them, for both continuous and discrete PINNs. For example, the law of mass action and the Arrhenius law are hard constraints in all the literature that considers them introduced in this paper, because they are explicitly embedded in the intermediate steps of the method rather than constrained by the loss. Hard constraints have both advantages and disadvantages. Imposing hard constraints on certain physical laws will alleviate the burden of loss optimization. For example, enforcing species conservation can reduce one equation of $Y$ that need to be solved. However, hard constraints can sometimes increase the difficulty of optimization, especially for IC/BCs, as they often involve hyperparameters and "hyperfunctions". In some cases, such as PDEs, complexity is also a disadvantage of hard constraints, which are less straightforward than soft constraints. The choice between soft constraints and hard constraints depends on many factors. By default, hard constraints should be considered first. Soft constraints can be used in the following situations: (1) when the constraint requirements are not strict; (2) when hard constraints are difficult to implement, such as for PDEs in continuous PINNs; (3) when hard constraints are extremely cumbersome to implement; (4) when hard constraints impair or greatly slow down the optimization and convergence, which requires a posteriori testing.
@@ 8.3 Opportunities for combustion and AI
  • Academic research. First, there are still many limitations that need to be further solved, as mentioned in Section 8.1, including large-domain long-time problems, stiffness caused by detailed reaction mechanisms, model-data-physics discrepancies, and reproducibility issues. Second, there are still many combustion scenarios that can be studied using PIML, such as multiphase combustion, pollutant control, ammonia combustion, microgravity combustion, explosions, and fires. Finally, there are still many tasks that can be achieved using PIML, such as improving combustion models, gaining insights from experimental data, and developing hybrid solvers.
@@
  • Integration of advanced AI methods. As introduced in this paper, various AI methods haven been used in PIML for combustion, including the basic MLP, KAN 283, NO, NODE, GAN, LSTM, AE, self-supervised learning, and transfer learning 304. However, there are still many advanced AI methods that have potential for application, such as DM, reinforcement learning, foundation models, and PDE discovery. Among these, DM has already been applied to fluid mechanics 169, 444, 445 and combustion 304, 446. Foundation models are a hot topic due to their notable applications in the NLP field, including LLMs such as GPT 25, Llama 447, and DeepSeek 448. In the SciML field, foundation models are a current key direction 89, 449. Some foundation models for PDEs have been developed, including the PDEformer 450, 451, UPS 452, Poseidon 453, PROSE-PDE 454, and PROSE-FD for fluid PDEs 455. Therefore, establishing foundation models for combustion is a feasible and promising direction 46, which also needs to be informed by combustion physics.
@@
  • Combustion-inspired AI. Physics-inspired AI paradigms are an important topic, in which AI methods are often proposed or improved by drawing inspiration from physical laws, phenomena, or principles. The inspiration can be achieved in many ways, including embedding physical laws, mimicking physical processes, and borrowing physical principles. Therefore, physics-inspired AI includes physics-informed AI, or it can be considered as physics-informed AI in a broad sense. Jiao et al. 456 have reviewed the AI methods that are inspired by classical mechanics, electromagnetism, statistical physics, and quantum mechanics. Additionally, two currently popular generative models are also inspired by physics: DM 163, 164 is inspired by nonequilibrium thermodynamics, while the continuity equation in the flow matching 457 originates from fluid mechanics. Therefore, it is natural to ask: can we develop combustion-inspired AI? In fact, CRNN can be regarded as one such effort 78, with its network architecture designed based on the law of mass action and Arrhenius law. However, given the abundance of combustion laws, phenomena, and principles, there are still many possibilities to explore, which may give rise to some revolutionary AI methods.
@@ 9 Conclusions
This paper provides a comprehensive review on PIML for combustion, systematically summarizing the fundamental principles, major studies, key advances, and available resources, while also providing constructive discussions and recommendations for researchers and engineers. The main conclusions are as follows: (1) PIML has been widely and successfully applied to combustion chemical kinetics, combustion reacting flows, and other combustion problems, addressing both forward and inverse problems. These applications demonstrate the growing attention, proven effectiveness, and significant potential of PIML in combustion research. (2) Numerous approaches exist for integrating combustion laws into ML. A primary focus of this review is to explain how such laws can be integrated into ML frameworks–either through soft or hard constraints, via loss functions or representation models, within coordinate-to-variable or field-to-field paradigms, and using either complete/detailed or reduced/simplified physics–highlighting the flexibility of PIML. (3) By incorporating combustion knowledge, the physical consistency and accuracy of ML models are always enhanced compared to purely data-driven ML. (4) PIML serves as a unified bridge between physics, models, and data, enabling both model-to-data simulation/reconstruction tasks and data-to-model modeling tasks under the guidance of physics This integration produces three key outcomes that represent or contain physics: enhanced data, improved physical models, and more reliable ML models. When discrepancies arise among physics, physical models, and data, PIML can address them through a range of strategies. (5) Challenges remain in terms of the inherent limitations of PIML, physics-model-data discrepancies, and reproducibility. Furthermore, many potential applications and research topics remain unexplored. Both of these points indicate promising directions for the future. In conclusion, PIML for combustion is a rapidly evolving field that calls for increased collaboration between combustion and AI researchers to address significant problems, develop improved strategies, and explore emerging opportunities.